# A Common Origin for Nano-Hz Gravitational Wave Background and Black Hole Merger Events

### v1

@@ Abstract
PBH accretion.— After their formation, PBHs can accrete surrounding mass and grow over time, while also losing mass through Hawking radiation. According to the generalized second law of thermodynamics, the accretion of PBH would dominate the radiation when the thermal bath temperature exceeds that of the BH, i.e., $T\gtrsim T_{\rm BH}$, where $T_{\rm BH}=1/(8\pi GM)$ represents the BH temperature. Zel’dovich and Novikov suggested that horizon-size PBHs could grow at a rate similar to the cosmological horizon before the end of the radiation era based on the Bondi-type accretion Zel’dovich and Novikov (1967). However, subsequent research by Ref. Carr and Hawking (1974) did not find a self-similar solution connecting a BH to an exact flat Friedmann background via a sound wave in a radiation-dominated Universe. The self-similar growth of PBHs may lead to a population with a mass around $10^{17}M_{\odot}$, but this has not been confirmed by observations Carr et al. (2010). Therefore, it is widely believed that horizon-size PBHs do not grow in the radiation era Carr et al. (2021).
@@
For redshift $z\lesssim z_{\rm dec}\simeq 130$, where the baryonic matter decouples from the CMB radiation, $\dot{m}$ remarkably increases due to the decrease of the relative velocity $v_{\rm rel}$ between PBHs and gas. Depending on the PBH mass, $\dot{m}$ peaks at redshift $z\sim 30$, with an accretion time scale $t_{\rm Salp}\sim 10^{8}$ yr, much larger than the age of the Universe $t_{\rm age}\sim 10^{7}$ yr. Therefore, the accretion of the PBH inside a virialized halo is significant during the redshift $10\lesssim z\lesssim 30$. For $z\lesssim 10$, the structure formation begins and the relative velocity increases to $v_{\rm rel}\gtrsim 10$ km/s, leading to a decrease of PBH accretion Ricotti et al. (2008). Other factors, such as reionization and global feedback, can also decrease the PBH accretion rate Ricotti et al. (2008); De Luca et al. (2020a). To address modeling uncertainties, we consider cut-off redshift values of $z_{\rm cut}=10$ and 7, as adopted in Refs. De Luca et al. (2020a) and Ricotti et al. (2008), respectively, below which the accretion is negligible.
In Fig. 1, the mass function of PBHs produced by the DW number density fluctuations is shown as the black curve at the benchmark point of $(T_{\rm ann},f_{w}(T_{\rm ann}))=(0.1~{}{\rm GeV},0.1)$, considering the temperature-dependent EOS Byrnes et al. (2018). We notice that the distribution peaks around $M\simeq 10~{}M_{\odot}$ and then sharply drops to zero due to DW annihilations. The PBH accretion effect with redshift cut-offs $z_{\rm cut}=10$ and 7 is represented by the red and blue curves, respectively. A smaller $z_{\rm cut}$ could lead to a larger accretion effect, leaving significant impacts on the interpretation of the LVK events with mass components $\gtrsim 10M_{\odot}$.
Figure 1: PBH mass function, assuming the DW number $N=3$.
GW phenomenology.— The annihilation of DW begins when $\Delta V_{\rm bias}\simeq\rho_{w}$, accompanied by intense GW radiation with a peak amplitude today Lu et al. (2024a, b)
$h^{2}\Omega_{\rm GW}^{\rm peak}(t_{0})\simeq 3.8\times 10^{-9}\left(\frac{f_{w}(T_{\rm ann})}{0.1}\right)^{2}$ (9)
and a peak frequency
$f_{\rm peak}(t_{0})\simeq 1.21\times 10^{-8}~{}{\rm Hz}\left(\frac{T_{\rm ann}}{0.1~{}{\rm GeV}}\right),$ (10)
where $g_{*}(T_{\rm ann})=10.75$ is adopted. Below $f_{\rm peak}$ the GW spectrum scales as $\propto(f/f_{\rm peak})^{3}$ from the causality requirement Caprini et al. (2009), whereas above $f_{\rm peak}$ simulations Hiramatsu et al. (2014) indicate a scaling behavior $\propto(f/f_{\rm peak})^{-1}$. We follow the model-independent procedure outlined in Ref. Lu et al. (2024b) to fit the PTA observations, with the results shown in Fig. 2. We observe that the interpretation of DW annihilation for the PTAs requires that the DW annihilation occur at the temperature $\sim 100$ MeV, around the QCD PT scale. We will show below that the benchmark point (the green star in Fig. 2) can also result in the correct PBH relic abundance from DW number density fluctuations by taking into account the PBH accretion effect. DWs may emit relativistic dark radiation and contribute to the effective number of neutrino species with $\Delta N_{\rm eff}\simeq f_{w}(T_{\rm ann})(p_{\rm DR}/0.16)(10.75/g_{*,s})^{1/3}$. The constraints on $\Delta N_{\rm eff}$ from current Planck Aghanim et al. (2020) and future Simons Ade et al. (2019) experiments are represented by the dashed and dotted lines, respectively.
Figure 2: The red and blue $1\sigma$ and $2\sigma$ regions represent the parameter space of a fit to the NANOGrav 15-year dataset Agazie et al. (2023) and IPTA-DR2 dataset Antoniadis et al. (2022), respectively.
To date, LVK has documented over one hundred BH merger events. While favoring BHs of astrophysical origin, some events may not align with stellar evolution predictions. For instance, GW190425 Abbott et al. (2020a), GW190814 Abbott et al. (2020b), and GW230529_181500 Abac et al. (2024) may have component masses exceeding neutron star or stellar BH upper/lower limits Clesse and Garcia-Bellido (2022); Chen et al. (2022), while the component masses of GW190521 fall within the mass gap $\sim(50-135)~{}M_{\odot}$ of the stellar-origin BHs Palmese and Conselice (2021), implying a primordial origin for this event De Luca et al. (2021). In Table 1, we present the average PBH mass $\langle M\rangle$ and the PBH energy fraction $f_{\rm PBH}\equiv\Omega_{\rm PBH}/\Omega_{\rm DM}$ from Eq. (6) using parameters near the benchmark point and find that $(\langle M\rangle,~{}f_{\rm PBH})\sim(10M_{\odot},~{}10^{-3})$. The PBH abundance is highly sensitive to the value of $f_{w}(T_{\rm ann})$ because of the exponential suppression in the Gaussian tail. As a reference, the Bayesian analysis of LVK O3 events in Ref. Andrés-Carcasona et al. (2024) favors the parameters $(16.8M_{\odot},~{}1.1\times 10^{-3})$ for the PBH origin.
@@
Table 1: The averaged mass $\langle M\rangle$ and the PBH density fraction $f_{\rm PBH}$ around the benchmark point, assuming $N=3$.
Note that the commonly used lognormal PBH mass distribution is strongly constrained by the LVK O3 events Andrés-Carcasona et al. (2024); Gow et al. (2020); Hall et al. (2020) due to the long positive tail in the high mass region, which predicts a high-mass binary merger rate larger than observations. Therefore, cutting off the high-mass tail of the lognormal distribution is expected to significantly enhance the Bayesian evidence Hall et al. (2020). We observe that the PBH mass distribution (6) has such a cut-off at high masses due to the DW annihilation.
Model implementation.— Let us now verify our scenario in a concrete particle model. Consider a scalar potential Chiang and Lu (2020)
@@
where the scalar $\Phi=(\phi+v_{\phi})e^{ia/v_{\phi}}/\sqrt{2}$ and $\mu_{3}$ is assumed to be real. The first term on the right-hand side of Eq. (11) respects a global $U(1)$ symmetry, which is softly broken to a $Z_{3}$ symmetry by the second term. The spontaneous breaking of $Z_{3}$ symmetry gives rise to the potential of the heavy axion $a$:
$V_{a}(a)=-\frac{1}{4}\mu_{3}v_{\phi}^{3}\cos\left(\frac{3a}{v_{\phi}}\right),$ (12)
with 3-fold degenerate vacua $a_{j}/v_{\phi}=2\pi j/3$, where $j=0,~{}1,~{}2$. DWs with a surface tension of $\sigma_{w}\simeq 8m_{a}f_{a}^{2}$ (where $m_{a}^{2}=9\mu_{3}v_{\phi}/4$ and $f_{a}=v_{\phi}/3$) are formed via the Kibble mechanism Kibble (1976) to separate the degenerate vacua.
The quantum anomaly of $Z_{3}$ under the color $SU(3)$ gauge group generates the topological interaction $\mathcal{L}\supset\frac{g_{s}^{2}a}{32\pi^{2}v_{\phi}}G^{\mu\nu}\widetilde{G}_{\mu\nu}$, where $G_{\mu\nu}$ is the gluon field tensor and $g_{s}$ is the strong gauge coupling. After QCD PT, the QCD instanton effects generate an effective potential for $a$
@@
where $\Lambda_{c}=(m_{u}\Lambda_{\rm QCD}^{3})^{1/4}\simeq 66.2$ MeV and $\vartheta$ is a strong CP-violating phase. The parameter $N_{c}$ depends on the number of quarks charged under the $Z_{3}$ symmetry, and we take $N_{c}=1$ to avoid color anomaly degenerate vacua. The potential (13) lifts the degeneracy of the 3-fold vacua by a bias energy $\Delta V_{\rm bias}^{1/4}=(1.5\Lambda_{c}^{4})^{1/4}\simeq 73.3$ MeV.
The condition $\rho_{w}\simeq\Delta V_{\rm bias}$ for DW annihilation gives
$\left(\frac{\sigma_{w}^{1/3}}{10^{5}~{}{\rm GeV}}\right)^{3}\simeq 39.6f_{w}(T_{\rm ann})\left(\frac{1}{\mathcal{A}}\right)\left(\frac{T_{\rm ann}}{0.1~{}{\rm GeV}}\right)^{2}$ (14)
@@
$\frac{\Delta V_{\rm bias}^{1/4}}{0.1~{}{\rm GeV}}\simeq 1.37(f_{w}(T_{\rm ann}))^{1/4}\left(\frac{T_{\rm ann}}{0.1~{}{\rm GeV}}\right).$ (15)
We then obtain $\sigma_{w}^{1/3}\simeq 1.5\times 10^{5}$ GeV and $\Delta V_{\rm bias}^{1/4}\simeq 77.0$ MeV for the benchmark model. Therefore, we have confirmed that the QCD instanton-induced potential (13) can indeed provide a correct bias energy for the decay of the degenerate vacua (12) in the benchmark model.
@@
Considering Eq. (4) and the given value $T_{\rm ann}\simeq 100$ MeV, it is possible to replicate the values of $\left(\langle M\rangle,~{}f_{\rm PBH}\right)$ in Tab. 1 for a model with arbitrary $N$ DWs by scaling $f_{w}(T_{\rm ann})$ with a factor of $\sqrt{N/3}$. In the clockwork axion model Choi and Im (2016); Kaplan and Rattazzi (2016), the value $f_{w}(T_{\rm ann})=0.2$ and the axion decay constant $f_{a}\simeq v_{\phi}3^{N}\sim 10^{11}$ GeV for $v_{\phi}\sim 10^{5}$ GeV and $N=12$ are obtained. While realizing our scenario, the axion with $f_{a}\sim v_{\phi}/\Lambda_{\rm QCD}^{2}\sim 10^{11}$ GeV provides a correct dark matter relic abundance through the misalignment mechanism Preskill et al. (1983); Abbott and Sikivie (1983); Lu et al. (2024a).
Conclusions.— We proposed that the number density fluctuations in a DW network can serve as a novel source for density perturbations, leading to the critical collapse and formation of PBHs, which is particularly significant near the annihilation of DWs by the bias potential induced by the QCD instanton effect during the QCD PT. PBHs with a mass distribution in the range of $(1-10)M_{\odot}$ can bridge the gap between $\sim(10-100)M_{\odot}$ for LVK by the accretion effects. By considering a benchmark model, we demonstrated that the mass distribution of PBHs resulting from DW Poisson fluctuations can explain the BH merger events, while the annihilation of DWs can generate the nano-Hz SGWB in PTAs. Additionally, we confirmed that the bias potential in the benchmark model can be accurately induced by the QCD instanton effect.
Future BH merger events by LVK beyond the stellar evolution are essential for testing our hypothesis, particularly through the analysis of the mass distribution and the spin of high-mass PBHs resulting from accretion De Luca et al. (2021). The high-luminosity LHC has the potential to probe a heavy axion with $f_{a}\sim(10^{3}-10^{5})$ GeV and mass in the range of $\sim(1-10^{3})$ GeV Baldenegro et al. (2018); Bauer et al. (2019); Ferreira et al. (2023). Additionally, the spontaneous breaking of $Z_{N}$ symmetry at the scale $v_{\phi}\sim 10^{5}$ GeV may lead to a first-order PT Chiang and Lu (2020), and the resulting GWs can be detectable for the LVK O3 run Chiang and Lu (2021).
Acknowledgments.— BQL is supported in part by Zhejiang Provincial Natural Science Foundation of China under Grant No. LQ23A050002. CWC is supported in part by the National Science and Technology Council under Grant No. NSTC-111-2112-M-002-018-MY3. TL is supported in part by the National Key Research and Development Program of China Grant No. 2020YFC2201504, by the Project No. 12275333 supported by the National Natural Science Foundation of China, by the Scientific Instrument Developing Project of the Chinese Academy of Sciences, Grant No. YJKYYQ20190049, and by the International Partnership Program of Chinese Academy of Sciences for Grand Challenges, Grant No. 112311KYSB20210012.

### v2

@@ Appendix A PBH accretion
After their formation, PBHs can accrete surrounding mass and grow over time while also losing mass through the Hawking radiation. According to the generalized second law of thermodynamics, the accretion of PBHs would dominate the radiation era when the thermal bath temperature exceeds that of the BHs, i.e., $T\gtrsim T_{\rm BH}$, where $T_{\rm BH}=1/(8\pi GM)$ represents the BH temperature. Zel’dovich and Novikov suggested that horizon-size PBHs could grow at a rate similar to the cosmological horizon before the end of the radiation era based on the Bondi-type accretion Zel’dovich and Novikov (1967). However, subsequent research in Ref. Carr and Hawking (1974) did not find a self-similar solution connecting a BH to an exact flat Friedmann background via a sound wave in a radiation-dominated Universe. The self-similar growth of PBHs may lead to a population with a mass around $10^{17}M_{\odot}$, but this has not been confirmed by observations Carr et al. (2010). Therefore, it is widely believed that horizon-size PBHs do not grow in the radiation era Carr et al. (2021).
@@
For redshift $z\lesssim z_{\rm dec}\simeq 130$, where the baryonic matter decouples from the CMB radiation, $\dot{m}$ remarkably increases due to the decrease of the relative velocity $v_{\rm rel}$ between PBHs and gas. Depending on the PBH mass, $\dot{m}$ peaks at redshift $z\sim 30$, with an accretion time scale of $t_{\rm Salp}\sim 10^{8}$ yr, much larger than the age of the Universe $t_{\rm age}\sim 10^{7}$ yr. Therefore, the accretion of the PBH inside a virialized halo is significant during the redshift $10\lesssim z\lesssim 30$. For $z\lesssim 10$, the structure formation begins and the relative velocity increases to $v_{\rm rel}\gtrsim 10$ km/s, leading to a decrease of PBH accretion Ricotti et al. (2008). Other factors, such as reionization and global feedback, can also decrease the PBH accretion rate Ricotti et al. (2008); De Luca et al. (2020a). To address modeling uncertainties, we consider cut-off redshift values of $z_{\rm cut}=10$ and 7, as adopted in Refs. De Luca et al. (2020a) and Ricotti et al. (2008), respectively, below which the accretion is negligible.
In Table 1, we show the average mass $\langle M\rangle$ and the PBH density fraction $f_{\rm PBH}$ around the benchmark point $(f_{w}(T_{\rm ann},T_{\rm ann})=(0.1,100{\rm MeV})$, assuming $N=3$. We observe that the accretion can be significant for PBH mass $M\gtrsim 10~M_{\odot}$. Therefore, a smaller $z_{\rm cut}$ could lead to a larger accretion effect, leaving significant impacts on the interpretation of the LIGO observations.
@@
Table 1: The average mass $\langle M\rangle$ and the PBH density fraction $f_{\rm PBH}$ around the benchmark point $(f_{w}(T_{\rm ann},T_{\rm ann})=(0.1,100{\rm MeV})$, assuming $N=3$.
Appendix B Model implementation
Let’s now construct a concrete particle physics model whose domain walls are annihilated by the QCD instanton-induced bias potential. Such a scenario could naturally generate the nano-Hz gravitational waves by the domain wall annihilation. Furthermore, the density perturbations from the Poisson fluctuations in the domain wall network could collapse to form PBHs that may account for the LIGO BH-BH merger events.
Consider a scalar potential Chiang and Lu (2020)
@@ Appendix B Model implementation
where the scalar field $\Phi$ is decomposed as
$\Phi=\frac{1}{\sqrt{2}}(\phi+v_{\phi})e^{ia/v_{\phi}},$
where $\phi$ and $a$ represent the radial and angular (axion) components, respectively, and $v_{\phi}$ is the vacuum expectation value (VEV) of $\Phi$. The parameter $\mu_{3}$ is assumed to be real. In Eq. (15), the first term on the right-hand side respects a global $U(1)$ symmetry, which is softly broken to a $Z_{3}$ symmetry by the second term.
To realize our framework, we consider the Kim-Shifman-Vainshtein-Zakharov (KSVZ) axion model Kim (1979); Shifman et al. (1980). This model introduces a heavy vector-like quark $Q$ and a scalar field $\Phi$. Under the standard model gauge group $SU(3)_{c}\times SU(2)_{L}\times U(1)_{Y}$, the charges of $Q$ and $\Phi$ are $(3,1,0)$ and $(1,1,0)$, respectively. The transformations of $Q$ and $\Phi$ under the global $Z_{3}$ symmetry are given by
$\Phi\to e^{i\alpha}\Phi,~~Q_{L}\to e^{i\alpha/2}Q_{L},~~{\rm and}~~Q_{R}\to e^{-i\alpha/2}Q_{R},$ (16)
where $\alpha=2\pi j/\mathcal{N}$ with $j=0,1,2,...,\mathcal{N}-1$ and $\mathcal{N}=3$ for the $Z_{3}$ symmetry. The spontaneous breakdown of $Z_{3}$ symmetry generates the potential of axion $a$:
$V_{a}(a)=-\frac{1}{4}\mu_{3}v_{\phi}^{3}\cos\left(\frac{3a}{v_{\phi}}\right).$ (17)
This potential exhibits three-fold degenerate vacua located at $a_{j}/v_{\phi}=2\pi j/3$ for j=0,1,2j=0,~1,~2. DWs with a surface tension of $\sigma_{w}\simeq 8m_{a}f_{a}^{2}$ (where $m_{a}^{2}=9\mu_{3}v_{\phi}/4$ and $f_{a}=v_{\phi}/3$) are formed via the Kibble mechanism Kibble (1976) to separate the degenerate vacua.
The Yukawa interaction for the heavy quark $Q$ with the scalar field $\Phi$ is expressed in the following Lagrangian:
$\mathcal{L}_{Y}=-(y_{Q}\Phi\bar{Q}_{L}Q_{R}+\text{h.c.}),$
where $y_{Q}$ is the Yukawa coupling, and $Q_{L}$ and $Q_{R}$ denote the left-handed and right-handed components of the heavy quark, respectively. When the scalar field $\Phi$ acquires a VEV $v_{\phi}$, the $Z_{3}$ symmetry is spontaneously broken, leading to a mass term for the quark:
$\mathcal{L}_{m}=-m_{Q}\bar{Q}_{L}Q_{R}e^{ia/v_{\phi}}+\text{h.c.},$
with $m_{Q}=\frac{y_{Q}v_{\phi}}{\sqrt{2}}$, and $a$ representing the axion field. To eliminate the phase $a$ from this mass term, one can implement a chiral transformation:
$Q_{L}\to e^{ia/2v_{\phi}}Q_{L},\quad Q_{R}\to e^{-ia/2v_{\phi}}Q_{R}.$
However, this transformation is anomalous under the QCD gauge group, resulting in a topological interaction between the axion and gluons, described by:
$\mathcal{L}\supset\frac{g_{s}^{2}a}{32\pi^{2}v_{\phi}}G^{\mu\nu}\widetilde{G}_{\mu\nu},$
where $G_{\mu\nu}$ is the gluon field strength tensor and $g_{s}$ indicates the strong coupling constant. After the QCD phase transition, this topological term generates an effective potential for the axion $a$ that is capable of breaking the $Z_{3}$ symmetry via QCD instanton effects:
@@
where $\Lambda_{c}=(m_{u}\Lambda_{\text{QCD}}^{3})^{1/4}\simeq 66.2$ MeV, and $\vartheta$ represents the strong CP-violating phase. The parameter $N_{c}$ relates to the number of quarks charged under the $Z_{3}$ symmetry and, to avoid complications linked to color anomaly degeneracies, we set $N_{c}=1$. The potential $V_{c}(a)$ serves to lift the three-fold vacuum degeneracy by introducing a bias energy:
$V_{\text{bias}}^{1/4}=(1.5\Lambda_{c}^{4})^{1/4}\simeq 73.3\text{ MeV}.$ (19)
This bias plays a crucial role in selecting a specific vacuum and significantly influences the dynamics of the axion field and its interactions with QCD.
Using the relations derived in Ref. Lu et al. (2024a)
$\left(\frac{\sigma_{w}^{1/3}}{10^{5}~{\rm GeV}}\right)^{3}=39.6f_{w}(T)\left(\frac{1}{\mathcal{A}}\right)\left(\frac{g_{*}(T)}{10.75}\right)^{1/2}\left(\frac{T}{0.1~{\rm GeV}}\right)^{2}$ (20)
@@
$\frac{V_{\rm bias}^{1/4}}{0.1~{\rm GeV}}=1.37(f_{w}(T_{\rm ann}))^{1/4}\left(\frac{T_{\rm ann}}{0.1~{\rm GeV}}\right)\left(\frac{g_{*}(T_{\rm ann})}{10.75}\right)^{1/4},$ (21)
we obtain $\sigma_{w}^{1/3}\simeq 1.5\times 10^{5}$ GeV and $V_{\rm bias}^{1/4}\simeq 77.0$ MeV for the benchmark model. Therefore, we have confirmed that the QCD instanton-induced potential (18) can provide correct bias energy for the decay of the degenerate vacua (17) in the benchmark model.